# Believing without Seeing: Quality Scores for Contextualizing Vision-Language Model Explanations

### v1

@@ Results.
Tables 1 and 2 compare three sets of explanation qualities: existing ones developed for text-only explanations (Simulatability, Informativeness, Plausibility), our proposed qualities (Visual Fidelity, Contrastiveness), and combinations of Visual Fidelity and Contrastiveness.
@@
We observe that on A-OKVQA, across all three VLMs, our proposed quality scoring functions achieve higher Discriminability scores and lower Expected Calibration Errors than existing scoring functions. We also see that all Discriminability scores are significant at $p<0.001$, indicating that both our quality scoring functions are capable of distinguishing accurate and inaccurate VLM predictions. Further, combining the Visual Fidelity and Contrastiveness quality scores results in even higher Discriminability scores and lower calibration errors, highlighting complementary utility. On VizWiz, we similarly find that Visual Fidelity is better calibrated than baseline quality scoring functions, and is better at distinguishing correct and incorrect predictions made by LLaVA.
Figure 2 shows calibration curves for the baseline qualities, our proposed qualities, and an average of Visual Fidelity and Contrastiveness, for explanations generated by Qwen2.5-VL-7B on the A-OKVQA task. We observe that Informativeness’s low calibration error is achieved by assigning high scores to all explanations. Our proposed quality scores achieve a lower calibration error while assigning a wider spread of scores across all evaluation samples.
@@ 4 Are VLM Explanation Quality Scores Helpful to Users?
Calibration alone does not necessarily translate to downstream utility for users in real-world scenarios Vodrahalli et al. (2022). To what extent can providing $\mathrm{VF}$ and $\mathrm{Contr.}$ scores alongside VLM explanations help users more accurately decide when to believe VLM predictions?
@@ 4.1 User Study Setup
We simulate a real-world setting of a user relying on a VLM assistant when the user cannot directly view the VLM’s visual input. Human participants assess the accuracy of VLM predictions.
Figure 3: Our study interface where users are shown the visual reasoning question, VLM prediction and explanation, and optionally one or more quality scores (using simplified language descriptions of both qualities). Here, the user believes that the VLM prediction is incorrect.
@@
We conduct a between-conditions user study, where users makes predictions about VLM correctness based on either none or a subset of the explanation quality scores. We conduct user studies on a subset of 100 questions each from the A-OKVQA and VizWiz tasks, with each subset containing 50 correct and 50 incorrect predictions by LLaVA-v1.5-7B. We select these subsets to approximately match the ECE of $\mathrm{VF}$ and $\mathrm{Contr.}$ scores (if available) with the full evaluation set to maintain a similar distribution of quality scores in our subset.
In a given setting (either none or some of the quality scores are presented to users), each question is annotated by three participants, yielding 300 annotations. Each participant completes 10 questions, also balanced to contain five instances where the model is correct and five where it is incorrect; participants are not informed of this balancing. Participants are randomly assigned to one of the study conditions (control or one of the treatment variants) and do not participate in more than one condition. In total, we collect 300 human annotations per dataset and condition (3 users × 100 questions), capturing user reliance behavior under varying explanation and quality signal conditions.
@@
Since participants are asked to judge whether the VLM’s prediction is correct or incorrect, we first exclude responses marked as “unsure.” Following prior work on evaluating human-AI collaboration Joshi et al. (2023); Ma et al. (2024); Srinivasan & Thomason (2025); Wiegreffe et al. (2022a), we then calculate User Accuracy over the remaining responses, and the degree of appropriate reliance using two metrics: Over-Reliance and Under-Reliance. Over-Reliance represents the fraction of interactions where the user believed the VLM prediction was correct, when in fact it was incorrect. Similarly, Under-Reliance is the fraction of interactions where the user mistakenly believed the VLM was incorrect. For both reliance metrics, lower values are preferred.
@@ Results.
Presenting explanation quality scores, either individually or in combination, consistently improves user performance across all three metrics (Figure 4). Relative to the control explanation-only condition and the Random Score baseline, all treatment conditions lead to higher user accuracy. On A-OKVQA, the most pronounced improvements are observed when showing the product of Visual Fidelity and Contrastiveness, which results in an 11.1% increase in user accuracy, a 15.4% reduction in over-reliance, and a 5.3% reduction in under-reliance. These gains suggest that quality signals help users not only better identify correct model predictions, but also avoid being misled by incorrect ones.
Other treatment variants also show meaningful improvements. Visual Fidelity and Contrastiveness alone both reduce over-reliance by approximately 2–4%, and displaying both scores side-by-side further improves user outcomes. On VizWiz, the trends are similar; addition of quality scores leads to improvements in both accuracy and over-reliance. Together, these findings highlight the utility of explanation quality scores in helping users make more informed and calibrated decisions.
Figure 5: Ablation on presentation types on A-OKVQA dataset: holding the numeric signal fixed (Prod($\mathrm{VF}$, $\mathrm{Contr.}$), $\mathrm{VF}$, or $\mathrm{Contr.}$), we vary the on-screen label (shown as VF, shown as Contr, or simple confidence). User Accuracy ($\uparrow$) and Over-Reliance ($\downarrow$) are effectively unchanged by naming/framing.
@@ 4.3 RQ2: Which aspects of the presentation affect user decision accuracy?
Building on RQ1, we explore which presentation choices are significant. We explore two dimensions:
(i) Labeling only (RQ2a): In this scenario, we display the numeric value, while the name displayed on the screen changes (e.g., it could be shown as $\mathrm{VF}$, $\mathrm{Contr.}$, or a generic “AI confidence”).
(ii) Format/content (RQ2b): Here, a scalar score is compared to a descriptive summary that presents evidence (e.g., verified visual claims for $\mathrm{VF}$ or plausible alternatives for $\mathrm{Contr.}$).
4.3.1 RQ2a: Do names and framings of a score affect user reliance?
One concern is whether our observed gains stem from what the score communicates versus how it is labeled to users. In A-OKVQA, using the same underlying score distributions (Prod($\mathrm{VF}$, $\mathrm{Contr.}$), $\mathrm{VF}$, and $\mathrm{Contr.}$), we permuted their on-screen presentation: (i) shown under the correct name, (ii) mislabeled as the other qualities or shown as a generic “AI confidence” without mentioning any quality.
@@ Results.
4.4 RQ2b: Do descriptive expressions of quality scores affect user decision accuracy?
Figure 7: Here, the study interface shows the participant descriptive text versions of the Visual Fidelity and Contrastiveness qualities.
Our proposed quality scores are interpretable. For Visual Fidelity, we can observe which visual details in the explanation were and were not verified by the verifier VLM. For Contrastiveness, we can examine which alternative answers were also assigned a high entailment probability. We compare the effect of showing explanation qualities as numeric scores versus interpretable text descriptions on user decision making. Figure 7 shows an example of our proposed qualities as text descriptions.
@@
As shown in Figure 6, we observe that descriptive formats perform comparably to, and in some cases slightly better than, their numeric counterparts. While showing both quality scores and for $\mathrm{VF}$ individually, the descriptive versions lead to improved user accuracy and reduced over-reliance. However, descriptive treatments hinder performance for $\mathrm{Contr.}$, leading to reduced accuracy. These findings suggest that while numeric scores offer a compact signal, descriptive formats may help reduce over-reliance by making the explanation evaluation process more transparent and interpretable.
@@
5 Related Work
@@
Prior work on evaluating explanation quality has largely focused on text-only domains, proposing metrics that assess informativeness Chen et al. (2023), prediction support Wiegreffe et al. (2022b); Hase et al. (2020), acceptability or helpfulness to users Wiegreffe et al. (2022a), and consistency across reasoning chains Golovneva et al. (2023); Prasad et al. (2023). Our work goes beyond textual plausibility by introducing two novel explanation qualities tailored for vision-language settings. Visual Fidelity has been explored for evaluating faithfulness of model-generated image captions Madhyastha et al. (2019); Lee et al. (2023). Our work builds on these ideas to do zero-shot evaluation of explanations.
@@
Most research on utility of natural language explanations for decision support has found that plausible-sounding explanations Jin et al. (2023) often mislead users into overrelying on inaccurate model predictions Joshi et al. (2023); Si et al. (2024b); Chaleshtori et al. (2024); Sieker et al. (2024). Our work adds to these findings by showing that evaluating and communicating explanation qualities to users can reduce overreliance.
6 Conclusions and Future Work
As VLMs are increasingly deployed in settings where users lack access to the model’s visual inputs, users must be able to discern when to rely on model predictions using secondary cues like explanations. However, providing explanations alone can mislead users into believing the model even when it is incorrect. We propose evaluating two unexplored, complementary explanation qualities: Visual Fidelity and Contrastiveness, and also introduce scoring functions for measuring both qualities. We find that our proposed quality scoring functions are well calibrated with model correctness, compared to existing notions of explanation quality. Through several user studies, we also demonstrate real-world utility of these scores, as users presented with these quality scores alongside VLM predictions and explanations are able to improve their task accuracy and reduce over-reliance on the VLM.
Limitations.
A-OKVQA multiple-choice questions may not be representative of real-world visual queries. Further, the design of our Contrastiveness quality limits its application to tasks with a closed-set of possible answers, which is often not available in realistic settings. Finally, we only evaluate on English-language datasets, and conduct user studies with only fluent English speakers.
Future Work.
An important direction is to develop adaptive human-AI reliance strategies that learn when to present explanations, when to suppress them, and when to show only quality scores, depending on the context, task difficulty, or user uncertainty. Additionally, explanation quality scores could be used as training objectives to improve the generation of explanations. Finally, future works should study how explanation quality influences user trust over time in order to better understand how users adapt to model signals and how trust dynamics evolve over repeated human-AI interactions.
Acknowledgments
Appendix A Code and Data Availability
@@
Appendix B Prompts used for Answers and Rationales Generation
We adopt a two‐step, post-hoc justification strategy ($IQ\to aE$) inspired by Wiegreffe et al. (2022b). In the first step, we prompt the model to produce only its answer prediction, without any accompanying rationale. In the second step, we present that predicted answer back to the model and ask it to “Please explain the reasoning behind your answer.”
@@ Appendix B Prompts used for Answers and Rationales Generation
Appendix C Existing Text-only Qualities Implementations
C.1 Simulatability
@@
Next, to transform the QA pair into an NLI task, we prompt GPT-4o-mini to convert the original question and its predicted answer into a single declarative hypothesis sentence $H$. Table˜4 shows the model and prompt used to generate $H$. We then feed the masked rationale $E_{masked}$ as the premise and $H$ as the hypothesis into the soumyasanyal/nli-entailment-verifier-xxl model on Hugging Face to obtain an entailment probability $p_{entail}$ as its Simulatability score.
@@ C.2 Informativeness
Table˜5 shows the model configuration and prompt used to evaluate informativeness of a rationale. After extracting the individual information pieces, we check the size of the resulting list; if it is non–zero, we deem the rationale to be informative.
Appendix D Visual Fidelity
@@ 5. Is the person not wearing sunglasses?
Table 7: Model configuration and prompt used to verify the visual questions generated from Table˜6.
@@
Table˜6 refers to the model setting and prompt used for generating the verification questions, while Table˜7 refers to those used for verifying the questions.
Appendix E Example of Model Explanations with Different Qualities

### v2

@@ Results.
Table˜1 and Table˜2 compare three sets of explanation qualities: existing ones developed for text-only explanations (Simulatability, Informativeness, Plausibility), our proposed qualities (Visual Fidelity, Contrastiveness), and combinations of proposed qualities.
@@
We observe that across all datasets and models, our proposed quality scoring functions almost always achieve higher Discriminability scores and lower Expected Calibration Errors than existing scoring functions across all models. Most differences in Discriminability on these datasets are statistically significant ($p<0.001$), confirming that our metrics effectively distinguish accurate from inaccurate predictions. We also find that our explanation-based metrics significantly outperform standard logit-based uncertainty baselines on complex reasoning tasks (see Appendix˜H). On VizWiz, while LLaVA-v1.5-7B shows strong discriminability, GPT-4o and Qwen2.5-VL-7B exhibit lower or negative scores on all scoring functions; as discussed in Appendix˜I, this is primarily due to valid answers being penalized by the dataset’s strict ground truth. Crucially, combining the Visual Fidelity and Contrastiveness quality scores results in the best overall Discriminability and Calibration, highlighting their complementary utility.
Figure 2 shows calibration curves for the baseline qualities, our proposed qualities, and an average of Visual Fidelity and Contrastiveness, for explanations generated by Qwen2.5-VL-7B on the A-OKVQA task. We observe that Informativeness’s low calibration error is achieved by assigning high scores to all explanations, while our proposed quality scores achieve a lower calibration error while assigning a wider spread of scores.
Verifier robustness.
To test whether VF’s behavior depends on the specific verifier model, we replace the default GPT-4o with two smaller open verifiers (Gemma-3n-E4B and Qwen2.5-VL-7B). Section˜J.3 shows that all three yield comparable discriminability and calibration, indicating VF remains stable across verifier architectures.
@@ 4 Are VLM Explanation Quality Scores Helpful to Users?
Calibration alone does not necessarily translate to downstream utility for users in real-world scenarios Vodrahalli et al. (2022). We now investigate whether providing $\mathrm{VF}$ and $\mathrm{Contr.}$ scores alongside VLM explanations helps users more accurately decide when to believe VLM predictions.
@@ 4.1 User Study Setup
We design a user study simulating a real-world setting of a user relying on a VLM assistant where the user cannot directly view the VLM’s visual input. Study participants are tasked with assessing the accuracy of VLM predictions.
Figure 3: Our study interface where users are shown the visual reasoning question, VLM prediction and explanation, and optionally one or more quality scores (using simplified language descriptions). Here, based on the explanation’s quality scores, the user correctly believes that the VLM prediction is incorrect.
@@
We conducted a between-subjects study on a balanced subset of 100 questions each from A-OKVQA and VizWiz (total 300 annotations per condition). Participants were recruited via Prolific and paid \$2 base + performance bonuses (see Appendix˜K for full protocol and demographics).
@@
Since participants are asked to judge whether the VLM’s prediction is correct or incorrect, we first exclude responses marked as “unsure.” Following prior work on evaluating human-AI collaboration Joshi et al. (2023); Ma et al. (2024); Srinivasan and Thomason (2025); Wiegreffe et al. (2022a), we then calculate User Accuracy over the remaining responses, and the degree of appropriate reliance using two metrics: Over-Reliance and Under-Reliance. Over-Reliance represents the fraction of interactions where the user believed the VLM prediction was correct, when in fact it was incorrect. Similarly, Under-Reliance is the fraction of interactions where the user mistakenly believed the VLM was incorrect. For both reliance metrics, lower values are preferred.
@@ Results.
Presenting explanation quality scores, either individually or in combination, consistently improves user performance across all three metrics (Figure 4). Relative to the control explanation-only condition and the Random Score baseline, all treatment conditions improve user accuracy. On A-OKVQA, the most pronounced improvements are observed when showing the product of Visual Fidelity and Contrastiveness, which leads to a 11.1% increase in user accuracy, a 15.4% reduction in over-reliance, and a 5.3% reduction in under-reliance. These gains suggest that quality signals help users better identify correct model predictions and avoid being misled by incorrect ones.
Other treatment variants also show meaningful improvements. Visual Fidelity and Contrastiveness alone both reduce over-reliance by approximately 2–4%, and displaying both scores side-by-side further improves user outcomes. On VizWiz, the trends are similar; addition of quality scores leads to improvements in both accuracy and over-reliance. Together, these findings highlight the utility of explanation quality scores in helping users make more informed and calibrated decisions about when to trust VLM predictions.
Figure 5: Ablation on presentation types on A-OKVQA dataset: holding the numeric signal fixed (Prod($\mathrm{VF}$, $\mathrm{Contr.}$), $\mathrm{VF}$, or $\mathrm{Contr.}$), we vary the on-screen label (shown as VF, shown as Contr, or simple confidence).
@@ 4.3 RQ2: Which aspects of the presentation affect user decision accuracy?
Building on our findings in RQ1, we explore which presentation choices are significant. We explore two dimensions:
  1. (i).
    Score labeling (RQ2a): We examine whether presenting the same numeric value, but varying its labeling on the study screen (e.g., presenting the score as $\mathrm{VF}$, $\mathrm{Contr.}$, or a generic “AI confidence”) affects user reliance.
  2. (ii).
    Score descriptiveness (RQ2b): Here, we examine the effect of presenting a scalar quality score versus a descriptive summary that presents evidence (e.g., verified visual claims for $\mathrm{VF}$ or plausible alternatives for $\mathrm{Contr.}$).
4.3.1 RQ2a: Does framing of the quality score affect user reliance?
One concern is whether our observed gains stem from what the score communicates versus how it is labeled to users. For instance, presenting a quality score as the visual fidelity of the explanation may yield different user reliance behavior than if the same score is presented as a generic AI confidence in the explanation.
In this ablation study, using the same underlying score distributions for A-OKVQA predictions (Prod($\mathrm{VF}$, $\mathrm{Contr.}$), $\mathrm{VF}$, and $\mathrm{Contr.}$), we permuted their on-screen presentation: (i) shown under the correct name, (ii) mislabeled as the other qualities or shown as a generic “AI confidence” that does not specifically mention either explanation quality.
@@ Results.
Figure 6: The descriptive quality presentation. The question and explanation display are identical to Figure˜3, but numeric scores are replaced with text descriptions of the Visual Fidelity and Contrastiveness qualities.
@@
4.3.2 RQ2b: Does providing language descriptions of quality scores improve user decision accuracy?
Our proposed quality scores are interpretable. For Visual Fidelity, we can observe which visual details in the explanation were and were not verified by the verifier VLM. For Contrastiveness, we can examine which alternative answers were also assigned a high entailment probability. We compare the effect of showing explanation qualities as numeric scores versus interpretable text descriptions on user decision making. Figure 6 shows an example of descriptive presentations of our proposed qualities.
@@
As shown in Figure 7, we observe that descriptive formats perform comparably to, and in some cases slightly better than, their numeric counterparts. While showing both quality scores and for $\mathrm{VF}$ individually, the descriptive versions lead to improved user accuracy and reduced over-reliance. However, descriptive treatments hinder performance for $\mathrm{Contr.}$, leading to reduced accuracy. These findings suggest that while numeric scores offer a compact signal, descriptive formats may help reduce over-reliance by making the explanation evaluation more transparent.
@@
5 Conclusions and Future Directions
While explanations are thought to help users discern when to rely on VLM predictions, providing explanations alone can mislead users into believing the model even when it is incorrect. We propose evaluating two unexplored, complementary explanation qualities: Visual Fidelity and Contrastiveness, and also introduce scoring functions for measuring both qualities. We find that our proposed quality scoring functions are well calibrated with model correctness, compared to existing notions of explanation quality. Through several user studies, we demonstrate real-world utility of these scores, as presenting users with quality scores alongside explanations improves their task accuracy and reduces over-reliance on the VLM.
An important direction is to develop adaptive human-AI reliance strategies that learn when to present explanations and quality scores and when to suppress them, potentially utilizing bandit algorithms to automate system-side abstention Feng and Boyd-Graber (2022). Additionally, explanation quality scores could be used as training objectives to improve the generation of explanations. Finally, future works should study how explanation quality influences user trust over time.
Limitations
One limitation of our approach lies in the open-ended Contrastiveness evaluation. Its score quality is inherently partially bounded by the semantic quality of the generated hard negatives. While we validate that our generator produces challenging distractors in Appendix˜I, the metric remains sensitive to their plausibility; if distractors are too easily distinguished, the metric may overestimate confidence, whereas if they are semantically ambiguous, it may penalize valid reasoning. Thus, our metric’s resolution is coupled with the generator’s capability. Second, while we demonstrate effectiveness on A-OKVQA, its multiple-choice format may not strictly be representative of real-world, open-ended visual queries. We attempt to bridge this by evaluating on open-ended VizWiz examples and the diverse MMMU-Pro benchmark, but future work should explore even broader task distributions. Finally, we only evaluate on English-language datasets and conduct user studies with fluent English speakers. Future work should explore multilingual generalization.
Ethics Statement
This work was conducted with careful consideration of ethical implications. Human studies were conducted with informed consent through Prolific, with participants compensated at rates exceeding US minimum wage (\$2 base + performance bonuses, see Section˜K.3.2). Participants were native English speakers from the US with higher education credentials (Table˜18). We acknowledge this may limit generalizability across demographic groups.
While our work aims to improve accessibility for blind and low-vision users by calibrating trust in VLM outputs, we recognize potential risks: (1) users might over-rely on quality scores themselves, (2) errors in quality assessment could disproportionately affect vulnerable populations, and (3) current English-only implementation excludes non-English speakers. Future work should address these limitations through multilingual support and extensive testing with blind/low-vision communities.
The VizWiz dataset used contains questions from blind and low-vision users but is fully anonymized. In our human studies, the annotators were anonymized. We filtered NSFW content (Section˜K.2) and did not collect any additional personal information. All model outputs were manually reviewed for offensive content before presentation to study participants.
A-OKVQA and MMMU-Pro are released under Apache-2.0 license. VizWiz is released under a CC BY 4.0 license. GPT-4o is accessed via OPENAI API under their terms of service. LLaVA-v1.5-7B and Qwen2.5-VL-7B are both under Apache-2.0 license. soumyasanyal/nli-entailment-verifier-xxl model we used for the NLI task (see Section˜D.1) is under a CC BY 4.0 license.
Our use of A-OKVQA, VizWiz, and MMMU-Pro aligns with their intended research purposes for evaluating VQA systems. We use language models consistent with their intended purposes.
We used AI assistants (e.g., ChatGPT) for grammar checking and code debugging, but not for producing substantive scientific content or analysis. All model outputs were reviewed and verified by the authors.
Acknowledgments
Appendix A Related Work
@@
Prior work on evaluating explanation quality has largely focused on text-only domains, proposing metrics that assess informativeness Chen et al. (2023), prediction support Wiegreffe et al. (2022b); Hase et al. (2020), acceptability or helpfulness to users Wiegreffe et al. (2022a), and consistency across reasoning chains Golovneva et al. (2023); Prasad et al. (2023). Our work goes beyond textual plausibility by introducing two novel explanation qualities tailored for vision-language settings. Visual Fidelity has been explored for evaluating faithfulness of model-generated image captions Madhyastha et al. (2019); Lee et al. (2023). Our work builds on these ideas to do zero-shot evaluation of explanations. Separately, recent work has investigated the calibration of verbalized confidence from models directly Xuan et al. (2025), whereas we derive calibration signals from the explanation content itself.
@@
While some studies in open-domain QA suggest that explanations can help users detect model errors González et al. (2021), many researches on utility of natural language explanations for decision support have found that plausible-sounding explanations Jin et al. (2023) often mislead users into overrelying on inaccurate model predictions Joshi et al. (2023); Si et al. (2024b); Chaleshtori et al. (2024); Sieker et al. (2024). Our work adds to these findings by showing that evaluating and communicating explanation qualities to users can reduce overreliance.
Appendix B Code and Data Availability
@@
Appendix C Prompts used for Answers and Rationales Generation
@@
Dataset User Prompt
MMMU-Pro Answer the following multiple-choice question. The last line of your response should be of the following format: ‘Answer: \$LETTER’ (without quotes) where LETTER is one of the options. Think step by step before answering. Question: {question}
Choices: {choices_with_letters}.
Table 4: Prompts used in MMMU-Pro VQA. The QA only contains one step. The model generates an explanation followed by an answer.
We adopt a two‐step, post-hoc justification strategy ($IQ\to aE$) inspired by Wiegreffe et al. (2022b). In the first step, we prompt the model to produce only its answer prediction, without any accompanying rationale. In the second step, we present that predicted answer back to the model and ask it to “Please explain the reasoning behind your answer”.
For MMMU-Pro, due to its complexity, we instead use a Chain-of-Thought strategy ($IQ\to Ea$), prompting the model to generate the explanation before the answer to facilitate reasoning (Table˜4).
Appendix D Existing Text-only Qualities Implementations
We compare our proposed quality measures against three established, text-only explanation qualities, each of which also returns a score in the range $[0,1]$. Simulatability Hase et al. (2020) evaluates whether an explanation offers sufficient evidence to logically justify the model’s prediction. Informativeness Chen et al. (2023) evaluates whether the explanation introduces new information to justify a prediction beyond just re-stating the question and prediction. Commonsense Plausibility Liu et al. (2023b) evaluates whether the explanation is in accordance with commonsense knowledge of everyday situations.
D.1 Simulatability
@@
Next, to transform the QA pair into an NLI task, we prompt GPT-4o-mini to convert the original question and its predicted answer into a single declarative hypothesis sentence $H$. Table˜5 shows the model and prompt used to generate $H$. We then feed the masked rationale $E_{masked}$ as the premise and $H$ as the hypothesis into the soumyasanyal/nli-entailment-verifier-xxl model on Hugging Face to obtain an entailment probability $p_{entail}$ as its Simulatability score.
@@ D.2 Informativeness
Table˜6 shows the model configuration and prompt used to evaluate informativeness of a rationale. After extracting the individual information pieces, we check the size of the resulting list; if it is non–zero, we deem the rationale to be informative.
D.3 Commonsense Plausibility
To evaluate the plausibility of the model’s explanations against commonsense knowledge, we utilize the Vera model Liu et al. (2023b) directly, without modification.
Appendix E Visual Fidelity
@@ 5. Is the person not wearing sunglasses?
Table 8: Model configuration and prompt used to verify the visual questions generated from Table˜7.
@@
Table˜7 refers to the model setting and prompt used for generating the verification questions, while Table˜8 refers to those used for verifying the questions.
Appendix F Hard Negatives Generations
To enable Contrastiveness evaluation on open-ended datasets like VizWiz (which lack fixed answer choices), we dynamically generate “Hard Negatives” to serve as plausible distractors. We use GPT-4o as the generator, providing it with the question, the ground-truth correct answer(s), and the image context. The model is instructed to create distractors that are incorrectly but visually plausible, explicitly avoiding simple synonyms. This aligns with prior work on automated distractor generation Lu et al. (2022); however, unlike previous learning-based methods, we employ a training-free approach using an off-the-shelf LLM. Table˜9 shows the exact prompt used.
Hard Negatives Generation Prompt
Given the following question, its correct answer(s), and the accompanying image, generate exactly THREE (3) different "Hard Negative" answers.
Context:
Question: {question}
Correct Answer(s): {valid_answers_str}
Your Goal: Generate 3 "Hard Negative" distractors.
A "Hard Negative" must be:
1. Incorrect: It must NOT be a synonym or a close grammatical variation of the correct answer(s).
2. Visually Plausible: Ideally, it refers to something else visible in the image or common in this context.
3. Challenging: It should be a reasonable guess for a Model that didn’t pay perfect attention effectively.
Constraints:
- NO Synonyms: If the answer is "bike", do NOT generate "bicycle".
- Format: Provide exactly THREE (3) comma-separated phrases. No numbering, no extra text.
- Length: Short, concise (1-3 words).
Table 9: System prompt used to generate hard negatives for Contrastiveness evaluation on open-ended datasets (VizWiz).
Appendix G Validation of Hard Negatives Strategy
To validate our strategy of dynamically generating hard negatives for open-set evaluation (as used in VizWiz), we conducted an ablation study on the A-OKVQA dataset. A-OKVQA typically provides 4 ground-truth answer choices (Closed-Set). We stripped these choices and instead generated hard negatives using the same pipeline applied to VizWiz, effectively creating a “Reconstructed Closed-Set” version of A-OKVQA.
Model Metric Closed-Set Reconstructed Closed-Set Delta
(Original) (Generated Hard Negatives)
GPT-4o Disc ($\uparrow$) $^{\phantom{***}}0.248^{***}$ $^{\phantom{*}}0.111^{*}$ 0.137-0.137
ECE ($\downarrow$) $0.210$ $0.312$ +0.102+0.102
Qwen2.5-VL-7B Disc ($\uparrow$) $^{\phantom{***}}0.283^{***}$ $^{\phantom{***}}0.158^{***}$ 0.125-0.125
ECE ($\downarrow$) $0.136$ $0.235$ +0.099+0.099
LLaVA-v1.5-7B Disc ($\uparrow$) $^{\phantom{***}}0.243^{***}$ $^{\phantom{***}}0.112^{***}$ 0.131-0.131
ECE ($\downarrow$) $0.176$ $0.233$ +0.057+0.057
Table 10: Ablation study on A-OKVQA comparing standard Closed-Set evaluation (using ground-truth options) vs. Reconstructed Closed-Set evaluation (using generated hard negatives) for the Contrastiveness quality. The generated negatives significantly increase task difficulty.
Table˜10 compares the performance of our Contrastiveness quality on the original Closed-Set vs. the Reconstructed Closed-Set. We observe a consistent drop in Discriminability ($\sim 0.13$) and a rise in ECE ($\sim 0.08$) across all models. This indicates that the generated hard negatives successfully make the task more challenging, likely because the generated distractors are semantically closer to the correct answer than the original multiple-choice options. However, crucially, the discriminability scores remain positive and statistically significant (e.g., $0.158^{***}$ for Qwen2.5-VL-7B), confirming that our metric remains valid and robust even in this harder, constructed setting.
To further understand why the generated negatives increase difficulty, we analyzed the semantic proximity of the distractors to the correct answer using all-mpnet-base-v2. Across the 500-sample set, the mean cosine similarity for Original Negatives was $0.424$, whereas for Generated Negatives it rose to $0.468$ ($\Delta+0.044$).
Image Question Correct Original Negatives (Sim) Generated Negatives (Sim)
[Uncaptioned image] What kind of coating has been used? frosting polish, paint, varnish
($S_{\text{sim}}=0.27$)
buttercream, whipped cream, glaze
($S_{\text{sim}}=0.58$)
[Uncaptioned image] In what type of location are they playing with the body board? room beach, park, store
($S_{\text{sim}}=0.28$)
living room, hallway, garage
($S_{\text{sim}}=0.58$)
Table 11: Comparison of Original (Closed-Set) vs. Generated (Reconstructed Closed-Set) distractors for matching A-OKVQA questions. The generated negatives are often semantically closer to the correct answer (higher similarity $S_{\text{sim}}$), making the discrimination task harder.
While A-OKVQA provides a rigorous baseline with plausible distractors, our Reconstructed Closed-Set method produces even more semantically proximal candidates. As shown in Table˜11, the generated options (e.g., “buttercream” vs “frosting”) often require finer-grained visual discrimination than original distractors (e.g., “paint” vs “frosting”). This suggests that LLM-generated hard negatives serve as a harder mode for evaluation, explaining the lower discriminability scores while verifying the metric’s robustness efficiently.
Appendix H Uncertainty Quantification Baselines
To situate our explanation quality metrics within broader uncertainty quantification literature, we compare them against a standard logit-based baseline. Specifically, for the LLaVA-v1.5-7B model (where we have access to output probabilities), we compute the softmax probability of the first token of predicted answer as a measure of model confidence.
A-OKVQA (Simpler) MMMU-Pro (Harder)
Metric Disc. $\uparrow$ ECE $\downarrow$ Disc. $\uparrow$ ECE $\downarrow$
[…]
VF $0.181^{***}$ $0.197$ $\mathbf{0.169}^{**\phantom{*}}$ $0.558$
Contr. $0.243^{***}$ $0.176$ $0.047^{\phantom{***}}$ $0.221$
Avg(VF, Contr.) $0.212^{***}$ $0.109$ $0.108^{**\phantom{*}}$ $0.360$
Prod(VF, Contr.) $\mathbf{0.320}^{***}$ $0.164$ $0.090^{**\phantom{*}}$ $\mathbf{0.144}$
Table 12: Comparison of our explanation quality metrics against a Logit-based Baseline (softmax probability of the predicted answer’s first token) on LLaVA-v1.5-7B. Significance is evaluated using an unpaired t-test: ∗∗∗ $p<0.001$, ∗∗ $p<0.01$, $p<0.05$.
Table˜12 presents the comparison on two contrasting datasets: A-OKVQA (simpler visual reasoning) and MMMU-Pro (complex, multi-discipline reasoning).
On A-OKVQA, the model’s intrinsic “intuition” (logits) is highly effective, achieving strong Discriminability ($0.252$) and excellent ECE ($0.046$). This suggests that for simpler tasks, raw confidence is a sufficient reliability indicator.
On MMMU-Pro, however, this signal breaks down. Logits yield poor discriminability ($0.055$, not significant) and high calibration error ($0.356$). In contrast, our explanation quality metrics—particularly the product of Visual Fidelity and Contrastiveness ($S_{\mathrm{Prod}}$)—remain robust, achieving statistically significant discriminability ($0.090^{**}$) and much better calibration ($0.144$).
These results indicate that logits (which measure answer confidence) and our metrics (which measure explanation reliability) are distinct signals. While logits suffice for simpler tasks, explanation quality offers a necessary and robust alternative for complex reasoning, particularly when model confidence becomes unreliable or inaccessible.